### v2

# Delayed takedown of illegal content on social media makes moderation ineffective

Abstract
Illegal content on social media poses significant societal harm and necessitates timely removal. However, the impact of the speed of content removal on prevalence, reach, and exposure to illegal content remains underexplored. This study examines the relationship with a systematic analysis of takedown delays using data from the EU Digital Services Act Transparency Database, covering five major platforms over a one-year period. We find substantial variation in takedown delay, with some content remaining online for weeks or even months. To evaluate how these delays affect the prevalence and reach of illegal content and exposure to it, we develop an agent-based model and calibrate it to empirical data. We simulate illegal content diffusion, revealing that rapid takedown (within hours) significantly reduces prevalence, reach, and exposure to illegal content, while longer delays fail to reduce its spread. Though the effect of delay may seem intuitive, our simulations quantify exactly how takedown speed shapes the spread of illegal content. Building on these results, we point to the benefits of faster content removal to effectively curb the spread of illegal content, while also considering the limitations of strict enforcement policies.
Introduction
Illegal and harmful content — including copyright violations, child sexual abuse material (CSAM), and incitement to violence or terrorism — poses serious harm and demands effective moderation efforts (1, 2). Platforms have developed a broad range of moderation tools and strategies that vary in severity. For some problematic content, such as misinformation and hate speech, interventions have been designed to promote critical user engagement, such as sensitive content labels (3, 4), accuracy reminders (5, 6, 7, 8), and friction mechanisms in content sharing (9, 10, 11). These “soft” interventions alter the environment to nudge user behaviors and are thus contingent on user motivations. Other techniques reduce the visibility of problematic content by limiting its searchability, recommendation, and engagement (12, 13, 14). Stricter interventions include content removal (takedown), account suspensions, and permanent bans (15, 16, 17). These approaches enable swift disruption of content virality and minimize exposure, but also raise concerns about censorship (18).
Given the trade-offs of different moderation approaches, previous modeling work has shown that combining interventions is an effective approach to leverage the strengths of each while mitigating their limitations (19). Nevertheless, for clear-cut cases of illegal content, removal or deactivation in concerned jurisdictions remains a legal demand for social media platforms. The importance of timely content removal has also been recognized in regulatory frameworks. Laws such as Germany’s Network Enforcement Act (NetzDG), Austria’s Communication Platform Act (KoPl-G), and France’s Loi Avia required platforms to remove “obviously illegal” content within 24 hours and other illegal content within seven days. These provisions reflected a clear regulatory emphasis on speed and have drawn criticism both for not being strict enough and for being too strict — potentially infringing on free expression (20, 21, 22, 23). One of the critical issues is the trade-off between moderation speed and accuracy: removing content very quickly can reduce its spread but may also lower precision, causing false-positive errors and legitimate concerns about over-moderation. The Digital Services Act (DSA) replaced precise deadlines with broader requirements: platforms must act “in a timely manner” (Art. 16(6)) and remove or disable access to illegal content “expeditiously” (Art. 6(1)). While such flexibility avoids rigid constraints, it also marks a shift away from legally enforceable time limits.
As legal requirements have shifted, the question of how takedown speed affects the prevalence, reach, and exposure of illegal content becomes highly relevant. However, thus far, policies enforcing takedown speed lack empirical grounding, making it difficult for policymakers to design and evaluate appropriate policies. This paper provides a comprehensive evaluation of the effects of moderation speed by proposing a model that measures how prevalence, exposure, and reach of illegal content change in response to takedown delay. In addition, we scrutinize available data on takedown delay of selected platforms and discuss the consequences and limitations of regulatory takedown deadlines.
Available research on the speed of content removal shows that delay in response reduces prevalence and exposure (24, 25). However, the existing studies largely focus on mis- and disinformation, such as misleading claims or conspiracy theories, which may not generalize to illegal content. Moreover, much of this work centers on emergencies, such as COVID-19 anti-vaccination narratives (26) or the January 6 Capitol Riot (25), during which platforms adopted exceptional measures. To inform legal frameworks, it is necessary to examine the standard operating procedures of platforms during routine, non-emergency periods. Studying the impact of content removal policies at scale is challenging, both because it would be ethically impermissible to experiment with harmful or illegal content, and because data on removals is often inaccessible due to privacy concerns. Currently, no studies systematically examine illegal content removal practices across platforms or quantify the effects of takedown delays.
Based on a year of data from the DSA Transparency Database (DSA-TDB) (27), we show that illegal content removal speed varies widely across platforms, ranging from several days to weeks, and in some cases, up to a year. By calibrating an agent-based model with empirically estimated takedown delays, we quantify the relationship between moderation speed and the reduction of illegal content spread. While longer delays predictably result in greater prevalence, reach and exposure of illegal content, our simulations reveal the precise, non-linear effect of takedown speed. The results indicate a steep increase in prevalence and exposure as takedown delay increases. Prompt removal of illegal content markedly reduces spread compared to systems with no or slow moderation. The impact of content removal becomes negligible as the delay extends beyond a couple of weeks — a timeframe that, according to our data, is typical for the takedown of illegal content on most platforms.
Our findings indicate that faster content removal can significantly reduce the spread of illegal material. However, current takedown delays often allow such content to circulate widely before action is taken. This suggests that regulatory frameworks like the DSA may benefit from stronger enforcement provisions aimed at promoting more timely and effective moderation. However, in the Discussion, we point to limitations of takedown deadlines and argue that any legal framework should balance the goal of timely removal with the need to ensure the quality of moderation and account for platform-level operational constraints.
Methods
Modeling social media diffusion
We use SimSoM (28), an agent-based model that mimics information diffusion on social networks such as Twitter/X, Mastodon, Bluesky, Threads, or Instagram. The platform is represented as a directed network, where nodes correspond to user accounts and edges denote follower relationships, which need not be reciprocal. As in real-world platforms, information circulates through messages that appear on user news feeds. At each time step, a user can post new original messages, reshare content from their feeds, or take no action. Messages represent information that could take the form of text, links, hashtags, images, or other media.
To select a message for resharing, an agent samples from the news feed, which contains distinct messages shared by their friends. Once shared by an agent, a message appears in the news feeds of their followers, allowing for propagation across the network. Thus, the exposure of the message depends on its author’s follower count — agents with more followers will achieve greater exposure. Capturing mechanisms of real-world ranking algorithms, agents prefer to reshare messages that are appealing, recent, and popular (29). The appeal of a message is a scalar value that captures the intrinsic attractiveness of messages. This could derive from properties such as novelty or emotional appeal. The distribution of message appeal captures a broad distribution of virality often seen in real-world information diffusion networks (28). The recency of a message decreases as time passes since it was first introduced to the agent’s news feed. The popularity of a message is the number times it was reshared by the agent’s friends.
Building on the base model, we incorporate additional features to examine the diffusion of illegal content. Specifically, each message includes a binary legality attribute indicating whether the content is legal or illegal. Different types of illegal content may be more or less appealing. For example, CSAM is criminally and morally condemned, therefore resharing is expected to be rare. More common forms of illegal content, such as copyrighted media, tend to be highly engaging. Lacking data on the correlation between legality and appeal, we make the simplifying assumption that the two are independent.
Figure 1: User activity distribution and probability of posting illegal content. (a) Complementary cumulative distribution (CCDF) of daily average Twitter user activity based on 81,633,118 posts by 37,882,999 accounts during nine days in January 2023 (orange line). The dashed line provides a scaling reference, with the slope 2.85-2.85 representing the exponent of the power law fit. (b) Illegal content posting distributions for low-risk (orange) and high-risk (purple) users, for mean illegal content probability $p=10^{-4}$ (top), $p=10^{-3}$ (middle), and $p=10^{-2}$ (bottom), based on a large random sample.
We also extend SimSoM to model the heterogeneous activity of social media users, where most agents have low activity levels, and only a few are hyperactive. We model this by assigning an activity level $a$ to each user at the beginning of the simulation. We draw this activity from a power law $P(a)\sim a^{\gamma}$, where we obtain the exponent $\gamma=2.85$ by fitting empirical measurements of average daily posts (Fig. 1a). If $a<1$, the user (re)shares a single message with probability $a$ at each time step; else the user (re)shares $\lfloor a\rfloor$ messages.
In the following sections, we outline our approach to examine the relationship between takedown delay and effectiveness. We first describe the modeling of illegal content diffusion and takedown, followed by an analysis of the DSA-TDB to estimate expected takedown times, and finally, we detail the experimental setup.
Modeling illegal content diffusion and removal
We assume that illegal content can be generated by any user, though only a small fraction $s_{H}$ of “high-risk accounts” (H) frequently share such content — for example, CSAM, scams, or terrorist propaganda. The rest of the network, “low-risk accounts” (L) may occasionally violate the law, such as sharing copyrighted content. The illegal content probability $p$ represents the aggregated probability across user groups, capturing the average likelihood that a user in the population posts illegal content. Specifically, the probability of posting illegal messages by both low- and high-risk accounts follows a beta distribution $P(i)\sim i^{\alpha-1}(1-i)^{\beta-1}$, where $\alpha>0$ and $\beta>0$ are parameters that express the mean $\frac{\alpha}{\alpha+\beta}$ and the skewness of the distribution for that group. Given the two group distributions, we can express the average illegal content probability across the full population as $p=s_{H}\frac{\alpha_{H}}{\alpha_{H}+\beta_{H}}+(1-s_{H})\frac{\alpha_{L}}{\alpha_{L}+\beta_{L}}$, where $s_{H}$ is the relative size of the high-risk groups, and the $\alpha$ and $\beta$ subscripts indicate the groups. Fig. 1b plots the probability distributions of posting illegal messages by each group for different values of $p$.
Since platforms typically do not provide access to original content, there is insufficient scrutiny to determine the legality of content and limited data on the share of illegal content (for an exception see Wagner et al. (30)). Hence, it is difficult to empirically calibrate the illegal content probability $p$ by estimating the percentage of illegal messages in the system. In our experiments, unless stated otherwise, we set $s_{H}=0.1$ to reflect a minority of high-risk accounts. We use the values of $\alpha_{L}=0.1$, $\beta_{L}=90$ for the low-risk group and $\alpha_{H}=3$, $\beta_{H}=30$ for the high-risk group, resulting in $p=10^{-2}$. We show that the results are robust to these parameter choices in the Robustness section.
To understand the dynamics of illegal content removal, let us consider the lifespan of such content on a platform. While some illegal content is immediately detected and removed using automated detection and decision systems, other is missed due to detection limitations. In such cases, content moderation is delayed until platforms either detect the content themselves later or receive external notification. Additional delays may occur when platforms assess the content against their community standards and national laws. If a violation is confirmed, either the content is removed globally or access is restricted in certain jurisdictions.
We model takedown delay by making the simplifying assumption that at any time step, there is a constant probability $p_{s}$ that a piece of illegal content is not detected and removed, and therefore survives. This yields an exponential survival curve for the delay $t$, $P(t)\sim 2^{-t/\tau}$, where the illegal content half-life $\tau$ is the time taken for half of the illegal content to be removed (median removal time since publication). We use $\tau$ to parameterize content takedown delay. For easier interpretation, we can alternatively define the expected time delay as $\tau/\ln 2$. Since $P(t)\approx p_{s}^{t}$ in the simulation, we set $p_{s}=2^{-1/\tau}$. When a message is removed, it is deleted from all user feeds. In the next section we derive a realistic range of $\tau$ values from empirical data for five major social media platforms.
DSA data analysis
Figure 2: Distributions (CCDF) of time delay in illegal content takedown on five platforms. Each plot reports empirical data from the DSA-TDB (green dots) along with an exponential fit (purple line). Top panels: Fits based on raw data with linear y-axes. Bottom panels: Fits based on a linear-log transformation of the data, highlighting extreme time delay values.
Under the DSA (Art. 15), platforms must publish Transparency Reports detailing “the median time needed for taking the action.” While platforms comply, reporting practices differ significantly, preventing thorough comparison among platforms. Moreover, median turnaround times in transparency reports only cover the time between notification and removal of content, not the time when problematic content circulates in the network before notification. Finally, DSA transparency reports do not indicate moderation delays specific to illegal content, and platforms disallow external access to verify data quality.
More detailed data, suited for our purpose, is available in the DSA Transparency Database, which compiles Statements of Reasons (SoRs) explaining single moderation decisions to affected users. These SoRs include the reasons for moderation, as well as information on dates of posting and moderation decisions. We therefore use this data source to estimate the empirical half-life ($\tau$) values. We collect SoRs from the DSA-TDB between January 1 and December 31, 2024. We capture relevant records using the strictest filter, where the platforms explicitly state that the moderation is due to illegal content. [^1] We select SoRs that have DECISION_GROUND_ILLEGAL_CONTENT in the decision_ground field. The database also includes the value DECISION_GROUND_INCOMPATIBLE_CONTENT for the same field. This can cover content that is illegal, which would be indicated in an optional boolean field incompatible_content_illegal. Since platforms do not use this reporting category consistently, we did not include it in our dataset for calibration. We consider five major social media platforms with a sufficient number of relevant SoRs: Facebook, Instagram, YouTube, TikTok, and Snapchat. We exclude LinkedIn and Pinterest due to a lack of SoRs related to illegal content and Twitter/X due to suspicious patterns revealed in its DSA data, as also highlighted in prior work (31). This results in 8,613 SoRs submitted by Facebook, 3,005 by Instagram, 408,400 by YouTube, 90,409 by TikTok, [^2] We excluded 2,949 observations with invalid input for the content creation date (2000-01-01). and 2,142 by Snapchat. The takedown delay is defined as the difference between the moderation date (application_date field) and the content posting date (content_date field).
Fig. 2 shows empirical survival time curves for the five social media platforms from DSA-TDB SoR records. By fitting exponential survival time curves $P(t)\sim 2^{-t/\tau}$ to this empirical data, we estimate that the half-life is $\tau\approx 21$ days for Facebook, 60 days for Instagram, 198 days for YouTube, 4 days for TikTok, and 94 days for Snapchat.
The exponential survival curves (Fig. 2 top) fit the data reasonably well for short delays, less for extended times. We can get an intuition for this by noting that the longer the survival of an illegal piece of content, the lower its visibility, and therefore the lower its likelihood to be reported and removed — $p_{s}$ increases instead of remaining constant. To better capture the tails of the data, we also estimate $\tau$ by fitting the linear-log transformed data to $\log_{2}P(t)\sim-t/\tau$. This procedure yields longer half-life estimates $\tau\approx 390$ days for Facebook, 340 days for Instagram, 786 days for YouTube, 249 days for TikTok, and 360 days for Snapchat (Fig. 2 bottom).
While none of the exponential survival curves perfectly fit the empirical takedown delay data, the short and long estimates from these different fits provide reasonable bounds for the half-life parameter. We therefore use them to determine a range of average delay for each platform.
Simulation framework and parameters
Based on the above analysis, we simulate various takedown scenarios by exploring a broad range of illegal content half-life values, $2^{-4}\leq\tau\leq 2^{9}$ days. Illegal content prevalence is defined as the average ratio of illegal content in circulation at time $t$: $I_{t}=\frac{1}{N}\sum_{i=1}^{N}f_{i,t}$ where $f_{i,t}$ is the fraction of illegal content in user $i$’s feed at time $t$ and $N$ is the number of agents.
The simulation ends once the system reaches a steady state, signaled by a convergence in the illegal content prevalence. To this end, we calculate an exponential moving average at time $t$ as $\bar{I}_{t}=\rho\bar{I}_{t-1}+(1-\rho)I_{t}$, where $\rho$ is a parameter regulating the importance of older values. We stop the simulation when the difference between two consecutive values of this moving average is smaller than a threshold, $|\bar{I}_{t}-\bar{I}_{t-1}|/\bar{I}_{t-1}<\epsilon$. The parameters $\rho=0.9$ and $\epsilon=0.0001$ used in the reported simulations were tested to ensure that the system’s illegal content prevalence stabilizes at the steady state.
For each set of parameters, we run 70 simulations without any content removal; these are averaged to obtain baseline metrics. We then run 70 simulations with content removal; illegal content prevalence and other quantities are averaged across these runs. All simulations start from random conditions.
Simulations are conducted on a follower network derived from empirical Twitter data. The network connects users obtained from a 10% random sample of public tweets (32). The dataset comprises users who posted links to news articles, with likely automated accounts excluded. The original follower network constructed from this dataset has 58,29658{,}296 nodes and 10,632,02310{,}632{,}023 edges. We reduce the size of this network to speed up the simulations in our main experiment (Effects of takedown delay section). We employ $k$-core decomposition to identify N=10,006N=10{,}006 nodes that constitute the $k=94$ core. Subsequently, a random selection of edges is removed to match the density of the original network. This procedure produces E=1,809,798E=1{,}809{,}798 edges, with each node having an average of roughly 180 friends or followers.
For the robustness analyses, we vary high-risk group size and illegal content probability (Robustness section). We therefore need a large follower network to run meaningful scenarios using low values of these parameters ($s_{H}=10^{-3}$ and $p=3\times 10^{-4}$). Therefore, we use the original network but remove a random set of edges to speed up the simulations. The average in/out-degree is thus reduced to 18 friends per node, resulting in a network with N=58,296N=58{,}296 nodes and E=1,063,202E=1{,}063{,}202 edges.
To test the robustness of our results with respect to different network structures, we create a third, synthetic network with features ubiquitous in real-world networks, namely, the presence of hubs and clustering (directed triads). To this end, we use a directed variant of the random-walk growth model (33). We configure the parameters of this model to generate a network comparable in size and average degree to the network used in the main experiment. In particular, we initialize the network with 181 fully connected nodes and assume new nodes have fixed out-degree $k_{out}=180$. New nodes are added by first randomly selecting one node to follow. For the remaining $k_{out}-1$ connections, each new node has an equal probability to follow a random friend of the initially chosen node or a randomly selected node in the network. This results in a network with N=10,006N=10{,}006 nodes and E=1,810,905E=1{,}810{,}905 edges.
Results
Takedown delay
Our analysis of the DSA-TDB for five major social media platforms reveals substantial variation in illegal content takedown delays — ranging from several days to weeks, and in some cases up to a year or more. TikTok has the fastest takedown speed, with most illegal content lasting approximately 6–359 days. YouTube has the the slowest content takedown time, where illegal content can take between 286 days and 3 years to be removed. The average illegal content life time for other platforms falls between these extremes, also ranging widely: approximately 30–563 days for Facebook, 87–491 days for Instagram, and 136–519 days for Snapchat. (See DSA data analysis for how these ranges are obtained.)
The observed heterogeneity in takedown delays may also reflect differences in the types of illegal content being moderated. Some content types, such as copyright violations, can often be verified quickly, while others may require more time for investigation.
Effects of takedown delay
We measure the effects of time delay in content takedown by the reduction of illegal content throughout the network, which can be quantified by multiple metrics. Prevalence measures the number of posts containing illegal content that are in the system at steady state (see Methods). The more illegal content is present in the system, the more harm it can potentially cause. However, a piece of illegal content may be removed from a user’s feed before the user sees it. Therefore we also measure harm more directly using two exposure metrics. Impressions measures the number of times illegal content is seen by users. Note that a single user can be exposed to many posts with illegal content, or even the same post multiple times. Reach is a complementary measure of exposure, capturing the number of unique users who see any amount of illegal content. For each of these measures, we calculate the reduction percentage relative to a baseline measure without any content removal. For example, the reduction in illegal content prevalence is calculated by $(\langle\bar{I}_{\text{baseline}}\rangle-\bar{I}_{\text{removal}})/\langle\bar{I}_{\text{baseline}}\rangle$, where $\bar{I}$ is the illegal content prevalence at the end of a simulation and $\langle\bar{I}_{\text{baseline}}\rangle$ is the baseline average.
Figure 3: Impact of takedown delay on illegal content. The reduction in the prevalence (pink), reach (green), and impressions (purple) of illegal content is plotted as a function of the expected takedown delay ($\tau/\ln 2$). Shading represents the 95% confidence intervals estimated using non-parametric bootstrapping. The vertical lines indicate the expected takedown delays obtained by fitting the DSA data to exponential curves for different platforms.
Fig. 3 plots the reduction in illegal content prevalence, impressions, and reach as a function of the expected takedown delay. Content removal has maximal impact under rapid takedowns. For expected delay below nine hours (corresponding to half-life $\tau<0.25$ days), removal results in approximately 95–100% reduction of illegal content across all measures. Moderation becomes less effective as the time delay increases, allowing illegal content to spread before being removed. Between approximately 9 hours–46 days ($0.25\leq\tau\leq 32$ days), the reduction in prevalence of illegal content drops from 95% to negligible levels. The empirical values measured from DSA-TDB data for TikTok and Facebook are within this interval. Those for Instagram, Snapchat, and YouTube correspond to longer delays, where removal has no significant effect on illegal content prevalence. The high ends of the delay ranges are even longer, as mentioned in the previous subsection. Although the overall trend is similar across the two exposure metrics, the reduction in impressions and reach drops even faster for delays of up to a few days. For example, when illegal content is removed within a day on average, reach and impressions are only reduced by about 60%.
In summary, these findings show that delays beyond a day significantly diminish the effect of content removal.
Robustness
The results presented above may depend on the values of $s_{H}$ (proportion of high-risk accounts), $p$ (illegal content probability), and the underlying follower network. We therefore investigate if the results remain valid when changing these parameters. For these robustness experiments, we use two different values of illegal content half-life, $\tau=2$ and $\tau=8$, and assign users to the high- and low-risk groups randomly in each run.
First, we run experiments to examine the effect of having smaller high-risk groups ($10^{-3}\leq s_{H}\leq 10^{-1}$), which may better reflect realistic scenarios. The main outcome variable, reduction in illegal content prevalence, shows no statistically significant differences (Fig. 4), suggesting that our findings are robust to changes in $s_{H}$.
Figure 4: Robustness of results with respect to different high-risk group sizes $s_{H}$ for $\tau=2$ (top) and $\tau=8$ (bottom) based on 20 simulations. The illegal content reduction does not vary significantly ($P\geq 0.72$ across pairs of experiments with different $s_{H}$ values, using Mann–Whitney U tests with Bonferroni correction).
Table 1: Data from the Future of Free Speech policy report and derived ratios of illegal content.
Disappeared/Total Legal/Disappeared Illegal Ratio
Germany (FB) 0.006 0.997 0.00002
Germany (YT) 0.115 0.989 0.001
France (FB) 0.012 0.921 0.0009
France (YT) 0.072 0.875 0.009
Sweden (FB) 0.005 0.946 0.0003
Sweden (YT) 0.041 0.946 0.002
Second, we wish to explore how the main result is affected by the volume of illegal content, which varies across platforms and countries. To estimate a plausible range for $p$, let us analyze the Future of Free Speech policy report (34). The report includes comment data collected from Facebook and YouTube across Germany, France, and Sweden during two weeks in June–July 2023. It also includes the fraction of comments that disappear ($\frac{\text{disappear}}{\text{total}}$) and the proportion of disappeared comments that are legal ($\frac{\text{legal}}{\text{disappear}}$). From this, we estimate the fraction of comments that are illegal: $\frac{\text{illegal}}{\text{total}}=\frac{\text{disappear}}{\text{total}}\times\frac{\text{illegal}}{\text{disappear}}=\frac{\text{disappear}}{\text{total}}\times(1-\frac{\text{legal}}{\text{disappear}})$. Table 1 shows that this apprach yields a range between 0.00002–0.009. We run experiments using these estimates (except for the smallest value, which would require a network with infeasibly large size, as discussed in Methods). In our main analysis we report the reduction in illegal content prevalence, $1-\frac{\bar{I}_{\text{removal}}}{\langle\bar{I}_{\text{baseline}}\rangle}$. To show that this ratio is independent of $p$, we plot the scaling relationship between $\bar{I}_{\text{removal}}$ and $\bar{I}_{\text{baseline}}$ in Fig. 5a for different $p$ values from Table 1. We observe only a weak sublinear relationship, $\bar{I}_{\text{removal}}\sim\bar{I}_{\text{baseline}}^{0.94}$, suggesting that the ratio is nearly independent of $p$. We also plot in Fig. 5b the reduction in the prevalence of illegal content for different values of $p$ from Table 1. We observe more noise for small $p$, but none of the differences between the values in our main results ($p=0.01$) and those for smaller $p$ values are statistically significant. These analyses suggest that our main results are robust with respect to the volume of illegal content in the system.
Figure 5: Robustness analysis for illegal content probability. (a) Scaling between illegal content prevalence in removal and baseline conditions. Each data point represents the proportion of illegal content in the system with and without removal. Results are based on 40 simulations using $\tau=2$. Each simulation is done with an empirical estimate of $p$ (vertical lines) for Facebook and YouTube users from Germany, France, and Sweden, derived from the data reported in Table 1. Estimates of $p$ are based on the proportion of disappeared content that is subsequently judged to be legal. The gray line represents a least-squares fit of the log-transformed data, yielding a slope of $0.94\pm 0.01$. (b) Illegal content reduction for different illegal content probabilities corresponding to the same countries and platforms, based on 10 simulations with $\tau=2$ (top) and $\tau=8$ (bottom). The result does not vary significantly ($P\geq 0.13$ across pairs of experiments with different $p$ values, using Mann–Whitney U tests with Bonferroni correction).
Last, we examine the effect of network structure by comparing experiments using the empirical network and a synthetic network (see Methods). There is no statistically significant difference in the reduction in illegal content prevalence between these two networks, indicating the robustness of our findings (Fig. 6).
Figure 6: Robustness analysis with respect to network structures based on 10 simulations with $\tau=2$ (left) and $\tau=8$ (right). The illegal content reduction does not vary significantly ($P=0.11$ and $P=0.22$, respectively, using Mann–Whitney U tests with Bonferroni correction).
Discussion
This paper reveals how delays in illegal content takedown impact its effectiveness. Using empirical takedown data for five social media platforms from the DSA-TDB, we find large variations in average takedown delays. While platforms like TikTok typically respond within a week, illegal content often remains online for over a month on platforms like Facebook, Instagram, Snapchat, and YouTube. Our agent-based model calibrated on this data shows that long delays can substantially diminish the impact of content removal. Although this relationship may appear intuitive, our analysis reveals non-linear dynamics between takedown speed and moderation effectiveness, including critical delay thresholds. Specifically, we show that removing illegal content within one day significantly reduces its prevalence and exposure, whereas prolonged delays allow illegal content to permeate user feeds — beyond 46 days, takedown becomes futile in curbing the spread of illegal content.
These findings have practical implications for social media moderation and regulation. By adopting more proactive takedown policies, platforms may reduce moderation delay and limit the spread of illegal content. Another approach to improving content moderation is to strengthen collaboration with trusted flaggers — individuals or organizations that identify and report content violating community guidelines or legal standards. With human flaggers being a limited resource (35), efforts should be directed towards prioritizing the most time-sensitive cases.
For legislation, the importance of timely moderation appears to support the argument for implementing tight illegal content removal deadlines, as imposed under NetzDG in Germany and KoPl-G in Austria prior to the DSA. However, the practical impact of such deadlines should be carefully assessed in regulatory debates for two reasons. First, regulatory takedown deadlines can only address the timespan between the notification of illegal content and its removal, not the period during which illegal content circulates in the network undetected. Therefore, the impact of tight deadlines on the spread of illegal content is inherently limited. In fact, most platforms already report fast reactions after notifications: a median time below 24 hours according to 2024 DSA Transparency Reports and removal of 80% of illegal content within 24 hours according to 2022 national Transparency Reports for Germany and Austria. These figures suggest that a large portion of the long delays measured from the DSA-TDB data occurs before illegal content is detected or flagged. To better inform policy, platforms should be required to report not only the timestamps for publication and removal of content, but also the detection/notification timestamps to the DSA Transparency Database. This would allow to identify how long content has circulated in the network undetected as well as the moderation delay in response to notifications.
Second, tight takedown deadlines may reduce the time necessary for properly assessing the (un)lawfulness of content, increasing the risk of wrong moderation decisions and thus lowering the quality of content moderation. Flagged content requires careful verification, including report reviews and appeal resolutions, before removal decisions can be made. Any regulatory framework must ensure effective moderation of illegal content while remaining practical for platforms, and mindful of potential adverse effects. For example, in France, the Loi Avia was struck down by the Constitutional Council (36) due to concerns about restrictions on freedom of expression (22, 23). These considerations suggest that platforms should remove immediately only obviously illegal content. For cases requiring further assessment, platforms should take time for thorough review. Until the assessment is completed, platforms may implement alternative interventions to reduce the circulation of potentially illegal content.
While our findings are robust, several limitations warrant consideration. The lack of empirical data on illegal content results in several assumptions about the content production process in our model. First, the model focuses on the spread of content already posted, without accounting for the user incentives or legal risks influencing decisions to post illegal material. We assume static probability distributions of sharing such content, though in reality, this behavior may evolve in response to moderation policies, potentially declining under stricter or faster moderation. Second, as mentioned in Methods, we assume uniform appeal and spreading potential across all types of content. If illegal content is generally more engaging, the model may underestimate the impact of delayed takedown, and vice versa. Lastly, we triangulate data from multiple sources, which may introduce inconsistencies. For instance, user activity is calibrated using global Twitter data, whereas the DSA database focuses on European users. This mismatch could lead to unknown biases in our calibration.
Our model of the moderation process also has limitations. The DSA-TDB data only captures the downstream outcomes of content moderation, neglecting the detection process. Our model thus treats all illegal content as equally detectable, which may overlook important variation in the detection of different content types. We also make the simplifying assumption that when illegal content is removed, it is deleted from all user feeds. In practice, however, some content may be illegal only in certain jurisdictions and access may be selectively disabled for users in these countries. Our model does not consider such a differentiation.
The database itself has limitations. It relies heavily on voluntary reporting by platforms, leading to inconsistent coverage and reporting quality both across and within platforms (31). This highlights a broader issue with self-reporting systems. Moderation decisions are shaped by platform-specific policies and guidelines, which may not always align with legal standards: content labeled as “illegal” may in fact be lawful but removed due to internal platform rules. The issue of false positives may have inflated the inferred probability of illegal content from the DSA-TDB.
Future research could extend our model by accounting for the fact that users often alter their posting patterns in response to moderation decisions (37, 38). One could also extend the model to compare the effects of takedown versus alternative types of interventions, as well as explore combinations of different intervention instruments. Finally, future models should consider the trade-offs between the harm caused by persistent illegal content and the unitended effects of moderation tools, e.g., the limitation of free speech caused by false-positive errors.
Author contributions
BT, SK, FM, FS, NJ, SG, and LL conceptualized the study. FM, NJ, SG, and LL acquired funding. BT, SK, and FM developed the methodology. FS and NJ analyzed the regulatory context and data sources. BT, SK, ES, EV, and GN performed formal analysis. BT and FM developed the software. FM provided computing resources. BT conducted the investigation. BT and SK visualized the data. BT, SK, FM, FS and NJ drafted the manuscript. All authors reviewed and edited the draft.
Data availability
Code and data to implement the model and reproduce results are available at github.com/osome-iu/simsom_removal.
Acknowledgements
We are grateful to Alessandro Flammini, Taekho You, Samuel Groesch, and Azza Bouleimen for helpful discussions; and to Nick Liu for the data collection used to estimate user activities. This work was supported in part by the Swiss National Science Foundation (Sinergia grant CRSII5_209250) and by the Knight Foundation.